OpenStudy (anonymous):

x+(x^(-1))/(-1)=x+1?

OpenStudy (johnweldon1993):

\[\large x + (\frac{x^{-1}}{-1}) = x+1\] We know x^-1 is just 1/x so we have \[\large x - \frac{1}{x} = x + 1\] We want the left side to look like the right side... we HAVE the x \[\large \color \red{x} - \frac{1}{x} = \color \red{x} + 1\] We just need to make \[\large - \frac{1}{x} = 1\] Well the only way THAT can happen...is if x = -1 Since \[\large - \frac{1}{-1} = 1\] So your answer would be x = -1

OpenStudy (anonymous):

\[\frac{ 4x-2 }{ x} =3\]